# RWKV-UNet: Improving UNet with Long-Range Cooperation for Effective Medical Image Segmentation

### v1

@@ Abstract
In recent years, there have been significant advancements in deep learning for medical image analysis, especially with convolutional neural networks (CNNs) and transformer models. However, CNNs face limitations in capturing long-range dependencies while transformers suffer high computational complexities. To address this, we propose RWKV-UNet, a novel model that integrates the RWKV (Receptance Weighted Key Value) structure into the U-Net architecture. This integration enhances the model’s ability to capture long-range dependencies and improve contextual understanding, which is crucial for accurate medical image segmentation. We build a strong encoder with developed inverted residual RWKV (IR-RWKV) blocks combining CNNs and RWKVs. We also propose a Cross-Channel Mix (CCM) module to improve skip connections with multi-scale feature fusion, achieving global channel information integration. Experiments on benchmark datasets, including Synapse, ACDC, BUSI, CVC-ClinicDB, CVC-ColonDB, Kvasir-SEG, ISIC 2017 and GLAS show that the RWKV-UNet achieves state-of-the-art performance on various types of medical image segmentation. Additionally, smaller variants, RWKV-UNet-S and RWKV-UNet-T, balance accuracy and computational efficiency, making them suitable for broader clinical applications.
1 Introduction
Computer-aided medical image analysis plays a crucial role in modern healthcare. Automated image segmentation is a critical component of this process, as it divides complex images into distinct regions, allowing specific tissues, organs, or lesions to be more clearly visible, improving diagnostic accuracy, and facilitating the development of personalized treatment plans. The rise of deep learning [28] has greatly advanced image segmentation by automatically learning intricate patterns within images with deep neural networks, improving the precision and efficiency of the analysis.
Current medical image segmentation approaches mainly rely on fully convolutional neural networks (FCNs) characterized by a U-shaped architecture. U-Net [46] is specifically designed for image segmentation tasks, particularly in the medical imaging domain. Its distinctive features include an encoder-decoder structure that enables effective downsampling and upsampling of images, along with skip connections. This design allows U-Net to capture both high-level semantic features and fine-grained details, resulting in improved segmentation accuracy. The architecture’s success has led to the development of numerous variants, such as Attention U-Net [40], U-Net++[61], U-Net3+ [22], which have further enhanced its performance and expanded its applicability.
While CNNs excel at capturing local features through fixed-size kernels, they often struggle with long-range dependencies needed for understanding complex structures. Transformers, with their self-attention mechanism, capture global context and intricate patterns across entire images. Combining these architectures allows U-Net to retain fine detail, as seen in models like TransUNet [9], MT-UNet [53], and HiFormer[21]. However, transformers have $O(N^{2})$ complexity in self-attention, which makes them computationally intensive for high-resolution scans. The trade-off between accuracy and efficiency is important in clinical scenarios, thus reducing computational complexity without significantly compromising accuracy is highly desirable.
Recently, sequence modeling methods based on linear attention mechanisms, such as Mamba  [17] and RWKV  [41], have attracted significant attention and been applied to computer vision  [62, 16] and medical image segmentation tasks  [47, 38, 33]. These models excel at capturing long-range dependencies with high accuracy, effectively preserving important contextual information across extensive image regions. In addition, their linear computational complexity greatly improves efficiency, making them suitable for large-scale or high-resolution images while reducing processing demands.
@@ 1 Introduction
Inspired by Vision-RWKV [47], we propose a novel RWKV-UNet, integrating convolutional layers with the spatial mix for medical image segmentation. The motivation behind this design is to take advantage of RWKV’s ability to capture long-range dependencies and combine it with the strengths of convolution localization, in order to achieve more precise and context-sensitive segmentation results. To evaluate our model’s effectiveness, we conducted experiments on benchmark datasets including Synapse  [27], ACDC  [5], BUSI  [1], CVC-ClinicDB  [4], CVC-ColonDB  [3], Kvasir-SEG  [25], ISIC 2017  [13] and GLAS  [49], where it achieves state-of-the-art (SOTA) performance, highlighting its potential as a robust tool for medical image analysis.
In summary, our contributions are threefold.
  • We propose a novel U-Net architecture based on a combination of RWKV and convolutional layers. An encoder based on the inverted residual RWKV (IR-RWKV) module is designed, effectively combining global and local information, with enhanced feature extraction capabilities through pre-training.
@@
  • A Cross-Channel Mix module is designed to improve the skip connections in U-Net, facilitating multi-scale information fusion and enhancing the model’s feature representation across different scales.
@@
  • Experiments conducted on datasets of different modalities show that the model achieves SOTA performance, demonstrating strong segmentation capabilities. We also propose relatively smaller models, RWKV-UNet-S and RWKV-UNet-T, which strike a balance between performance and computational efficiency, making them suitable for a wider range of scenarios.
2 Related Work
2.1 UNet Variants
@@
U-Net is a deep learning architecture designed for image segmentation, originally developed for biomedical images. Its symmetrical U-shaped structure consists of an encoder and a decoder, where the encoder progressively extracts features and reduces spatial dimensions, while the decoder gradually restores spatial information and generates high-resolution outputs. Skip connections are used to combine feature maps from the encoder layers with those from the decoder layers, ensuring the preservation of details. Variations in U-Net improve its performance through various means, including the use of deeper encoders such as ResNet  [59], the introduction of gating mechanisms to focus on important features  [40], improvements in skip connection formats  [61, 22, 64, 42] to recover more spatial details, as well as adaptations for 3D images [12, 39]. Some studies have explored combining MLP/KAN with UNet to achieve long-range dependencies  [50, 35, 30]. These improvements allow U-Net to demonstrate higher accuracy and adaptability in various image segmentation tasks.
Attention-based Improvements.
Attention-based U-Net has several advantages, primarily its ability to capture global context information, which enhances the model’s understanding of long-range dependencies. Furthermore, the self-attention mechanism of Transformers allows the model to focus more effectively on key information. Variants of attention-based U-Net include pure Transformer models  [19, 23, 2, 8], which rely entirely on self-attention mechanisms for feature extraction and image segmentation. Another category consists of hybrid architectures that combine Transformers with CNNs  [9, 7, 53, 57, 52, 55, 21, 26], which leverage the local feature extraction capabilities of CNNs along with the global information processing strengths of Transformers to better capture image details and structural information. Recently, Mamba  [17]’s success in NLP and vision tasks  [62] has drawn significant attention to SSM, which can be considered a form of linear attention. SSM-based UNet variants  [47, 38, 54, 34, 33] capture long-range dependencies effectively and offer linear computational complexity, enhancing accuracy and efficiency in image segmentation tasks.
2.2 Receptance Weighted Key Value (RWKV)
Receptance Weighted Key Value (RWKV)  [41] was initially developed for natural language processing tasks. It combines the linear complexity of RNNs with the parallel processing advantages of Transformers, resulting in excellent performance, constant memory usage, and consistent inference generation speed. Building on this, Vision-RWKV (VRWKV)  [16] introduces RWKV to computer vision tasks, achieving success in image classification, object detection, and semantic segmentation. It efficiently handles sparse inputs and robust global processing and allows for high-resolution image processing without the need for window operations. Related efforts have also emerged in medical image analysis tasks. Restore-RWKV  [56] is an RWKV-based model for medical image restoration, enhancing the efficiency and effectiveness of medical image restoration through the introduction of the linear complexity Re-WKV attention mechanism and the omnidirectional token shift (Omni-Shift) layer. BSBP-RWKV [60, 60] achieves superior accuracy in medical image segmentation by combining the noise reduction advantages of Perona-Malik diffusion with the efficient RWKV structure.
3 Methodology: RWKV-UNet
3.1 Overall Architecture
@@
Figure 1: Overall architecture of the proposed RWKV-UNet. (a) the encoder constructed by stacked IR blocks and stacked IR-RWKV blocks, (b) Cross-Channel Mix (CCM) Module, (c) the decoder constructed by point convolution layers and depthwise convolution layers,(d) the Inverted Residual (IR) block, (e) the Inverted Residual RWKV (IR-RWKV) block, (f) the architecture of the decoder block, containing a point-convolution layer and a $\times 9$ 9 DW-Conv layer, following a point operation layer and an upsampling operation.
@@
The overall architecture of the proposed RWKV-UNet is presented in Figure 1. RWKV-UNet consists of an encoder, a decoder and skip connections with a Cross-Channel Mix (CCM) Module.
3.2 Effective Encoder in RWKV-UNet
Inverted Residual RWKV (IR-RWKV) Block.
@@ 3.2 Effective Encoder in RWKV-UNet
We employ the spatial-mix module in Vision RWKV and Depth-Wise Convolution (DW-Conv) with a skip connection to combine global and local dependency, and trade-off model cost and accuracy. The process of expanding and reducing dimensions enhances feature representation by capturing richer details while preventing information bottlenecks. Given a feature map $X\in\mathbb{R}^{C_{\mathrm{in}}\times H\times W}$, the processing flow through the IR-RWKV module is as follows:
Normalize and project the input $X$ to an intermediate dimension $C_{mid}$ using a $\times 1$ 1 Convolution layer:
@@
Divide the feature map into patches of size $\times 1$ 1. Then the Spatial Mix in VRWKV is applied:
@@
Convert the feature sequence back into a 2D feature map $I_{4}=\text{Folding}(I_{3})\in\mathbb{R}^{C_{\mathrm{mid}}\times H\times W}$, and use a $\times 5$ 5 DW-Conv layer for local feature aggregation:
@@
Inverted Residual (IR) Block.
An IR block contains a point convolution layer, a DW-Conv layer and a point convolution layer with local and global residual skips, removing the spatial mix as well as the unfolding and folding processes from the IR-RWKV.
Effective Encoder with Stacked IRs and IR-RWKVs.
We construct the RWKV-UNet encoder by stacking the IR and IR-RWKV blocks. The encoder comprises a stem stage followed by four stages. The first blocks in stages \@slowromancapi@, \@slowromancapii@, \@slowromancapiii@ and \@slowromancapiv@ do not have residual connections and use a DW-Conv with a stride of 2 to achieve downsampling. After the first block, Stages \@slowromancapi@ and \@slowromancapii@ are composed of stacked IR blocks without Spatial Mix. In Stages \@slowromancapiii@ and \@slowromancapiv@, a series of IR-RWKV blocks follow the first block.
Scale Up.
As shown in Table 1, we implement three different sizes of encoders with different numbers of IR and IR-RWKV blocks in each stage (depths), different embedding dimensions and expansion ratios for $C_{\mathrm{mid}}$ : Encoder-Tiny (Enc-T), Encoder-Small (Enc-S), and Encoder-Base (Enc-B), allowing flexibility to scale up the model according to different resource constraints and application needs.
Pre-training Manner.
@@
Table 1: RWKV-UNet Encoder Configurations Across Stages
@@
Parameters 2.6M 9.4M 16.7M
FLOPs 1.9G 4.8G 9.0G
Table 2: Comparison results of models with different attention mechanisms on Synapse dataset.
@@
No Attention 18.06 77.05 04.55G
Multi-Head Self-attention [51] 16.95 77.60 14.89G
Focused Linear Attention [18] 18.45 75.86 11.69G
Bi-Mamba [62] 13.13 77.86 14.53G
Spatial-Mix [47] 14.03 78.68 11.22G
@@
We conduct experiments by replacing the spatial mix in the encoder of RWKV-UNet with other attention mechanisms, such as self-attention  [51], focused linear attention  [18] and Bi-Mamba used in Vision Mamba  [62], as well as by removing this module entirely. The total training epochs are 60. The results shown in Table 2 indicate that the model with Spatial-Mix achieves the best DSC with the lowest computational cost compared to other attention mechanisms, showcasing its efficiency and effectiveness. And all kinds of attention mechanisms can lead to improvement compared to pure CNN structure except focused linear attention.
3.3 Decoder Design
We design our CNN-based decoder block with a point-convolution layer and a $\times 9$ 9 DW-Conv layer, following a point operation layer and an upsampling operation. Consider the input feature map $X\in\mathbb{R}^{C_{\text{in }}\times H\times W}$; in depthwise convolution, the number of input channels is $C_{\mathrm{in}}$, and the second point convolution maps the $C_{\text{in }}$ channels to $C_{\text{out }}$. The computational complexity of a standard convolution is:
$H\times W\times C_{\text{in }}\times\left(C_{\text{out }}\times K^{2}\right).$ (6)
The computational complexity of pointwise convolution + depthwise convolution + pointwise convolution is:
$H\times W\times C_{\text{in }}\times\left(C_{\text{in }}+K^{2}+C_{\text{out }}\right),$ (7)
where $K$ is the kernel size. The latter method is more computationally efficient for the use of larger kernels, as the complexity is proportional to $C_{\text{in }}\times K^{2}$ rather than $C_{\text{in }}\times C_{\text{out }}\times K^{2}$.
3.4 Cross-Channel Mix for Multi-scale Fusion

### v2

@@ Abstract
In recent years, significant advancements have been made in deep learning for medical image segmentation, particularly with convolutional neural networks (CNNs) and transformer models. However, CNNs face limitations in capturing long-range dependencies, while transformers suffer from high computational complexity. To address this, we propose RWKV-UNet, a novel model that integrates the RWKV (Receptance Weighted Key Value) structure into the U-Net architecture. This integration enhances the model’s ability to capture long-range dependencies and to improve contextual understanding, which is crucial for accurate medical image segmentation. We build a strong encoder with developed Global-Local Spatial Perception (GLSP) blocks combining CNNs and RWKVs. We also propose a Cross-Channel Mix (CCM) module to improve skip connections with multi-scale feature fusion, achieving global channel information integration. Experiments on 11 benchmark datasets show that the RWKV-UNet achieves state-of-the-art performance on various types of medical image segmentation tasks. Additionally, smaller variants, RWKV-UNet-S and RWKV-UNet-T, balance accuracy and computational efficiency, making them suitable for broader clinical applications.
I Introduction
Computer-aided medical image analysis is crucial in modern healthcare, where deep learning–driven automated segmentation precisely delineates anatomical structures and lesions, thereby enhancing visualization, improving diagnostic accuracy, and supporting personalized treatment planning. U-Net [1] is a seminal segmentation architecture featuring an encoder-decoder structure with skip connections, allowing it to capture both high-level semantic context and fine-grained details, which has inspired numerous improved variants.
Medical images are typically high-resolution and exhibit complex anatomical structures with diverse lesion appearances, posing challenges for segmentation tasks that demand both fine local detail and coherent global context. Local features are essential for accurate boundary delineation and small lesion detection, whereas long-range dependencies ensure structural consistency and cross-slice contextual integration. CNN-based designs excel at local feature extraction through convolutional kernels but struggle to model long-range dependencies. Transformers-based [2, 3] UNet variants like [4, 5] address this via patch-based image processing and self-attention, improving segmentation accuracy, but the $O(N^{2})$ computational cost impacts efficiency, especially when processing high-resolution images.
Figure 1: Comparative analysis of CNN-, Transformer-, Mamba-, RWKV-, and hybrid-based segmentation models, highlighting their respective strengths and weaknesses.
Recent linear-attention models such as Mamba [6, 7] and RWKV [8, 9] have emerged as efficient alternatives, both achieving linear complexity while maintaining long-range modeling capability, and UNet variants based on these [10, 11, 12] have been developed for medical image segmentation. Mamba leverages state-space dynamics and selective scanning for efficient global information propagation, whereas RWKV introduces a gated recurrent mechanism that combines recurrence inductive bias with Transformer-style context aggregation. From a modeling perspective, RWKV’s gated accumulation enables more persistent and isotropic context propagation capability, yielding broader effective receptive fields and more stable long-range dependency modeling in medical image segmentation. Integrating CNNs with these global modeling mechanisms unifies precise local feature extraction and efficient global context understanding, addressing the inherent limitations of either CNNs or Transformers alone. Building on RWKV’s advantages, CNN–RWKV hybrids are expected to provide more balanced representations, offering a promising trade-off between segmentation accuracy and computational efficiency. Fig. 1 compares the strengths and weaknesses of each model.
@@ I Introduction
Based on the advantages discussed above, we propose a novel RWKV-UNet for medical image segmentation, explicitly integrating convolutional layers with the spatial mixing capabilities of RWKV. This design leverages RWKV’s strength in capturing long-range dependencies and directionally consistent context propagation while retaining the precise local feature extraction of convolution, aiming to produce more accurate and context-aware segmentation results. We evaluate our model on 11 benchmark datasets, where it achieves state-of-the-art (SOTA) performance, demonstrating its effectiveness and potential as a robust tool for medical image analysis.
In summary, our contributions are as follows:
  • We propose an innovative RWKV-based Global-Local Spatial Perception (GLSP) module to build encoders for medical image segmentation. This approach efficiently integrates global and local information, and its feature extraction capabilities are augmented via pre-training.
@@
  • We design a Cross-Channel Mix (CCM) module to improve the skip connections in U-Net, which facilitates the fusion of multi-scale information and enhances the feature representation across different scales.
@@
  • Our RWKV-UNet demonstrates superiority and efficiency, achieving SOTA segmentation performance on 11 datasets of different imaging modalities. Additionally, the relatively compact models, namely RWKV-UNet-S and RWKV-UNet-T, strike a balance between performance and computational efficiency, making them adaptable to a wider range of application scenarios.
II Related Work
II-A Architectures for Medical Image Segmentation
@@
U-Net [1] is a deep learning architecture for biomedical image segmentation, featuring a U-shaped encoder-decoder structure with skip connections that preserve spatial details while extracting and restoring features. Variations based on U-Net improve its performance through various means, which include using developed encoders [13, 14], gating mechanisms to focus on important features [15], improving skip connection formats [16, 17, 18] to recover more spatial details, as well as adaptations for 3D images [19, 20]. Some studies have explored combining MLP/KAN with UNet to capture long-range dependencies [21, 22, 23]. nnU-Net [24] stands out as a self-configuring framework that automatically adapts preprocessing, architecture, and training settings to different medical datasets, serving as a robust baseline in many clinical segmentation challenges. These improvements allow U-Net to demonstrate higher accuracy and adaptability in various image segmentation tasks.
Attention-based Improvements. Attention-based methods have demonstrated strong performance in visual tasks due to their ability to capture global context. The self-attention mechanism of Transformers allows models to focus on the most informative regions, which is crucial in medical image segmentation for precise delineation of anatomical structures and lesions. Pure Transformer U-Nets [4, 5, 25, 26, 27] rely entirely on self-attention, enabling effective long-range dependency modeling but often at high computational cost. Hybrid CNN-Transformer architectures [28, 29, 30, 31, 32] combine CNNs’ local feature extraction with Transformers’ global modeling, preserving fine structural details while capturing complex tissue and lesion relationships. More recently, linear-attention sequence models such as Mamba [6, 7] and SSM-based U-Net variants [10, 11, 33, 34, 35] efficiently capture long-range dependencies with linear complexity, offering a favorable balance between segmentation accuracy and computational efficiency. And there are also attention-based attenpt for segmentation in medical videos [36]. Despite the progress,existing parallel global-local fusion suffers from high computational cost and can interfere with fine feature representation, while some global modeling may be unnecessary in shallow layers. Additionally, existing approaches may struggle to capture long-range dependencies stably, as they lack recurrent inductive bias that support stable context propagation.
II-B Receptance Weighted Key Value (RWKV)
Receptance Weighted Key Value (RWKV) [8] models sequential dependencies via a weighted combination of past key-value pairs modulated by a learnable gate, avoiding the quadratic cost of traditional attention. Adapted to vision tasks as Vision-RWKV (VRWKV) [9], it efficiently captures long-range spatial dependencies in high-resolution images with linear complexity and reduced memory overhead, and has also been extended to broader visual domains [37, 38].
RWKV for Medical Imaging. Restore-RWKV [39] applies linear Re-WKV attention with an Omni-Shift layer for efficient multi-directional information propagation in image restoration. BSBP-RWKV [40] integrates RWKV with Perona-Malik diffusion to achieve high-precision segmentation, preserving structural and pathological details. RWKVMatch [41] leverages Vision-RWKV-based global attention, cross-fusion mechanisms, and elastic transformation integration to handle complex deformations in medical image registration. Zigzag RWKV-in-RWKV (Zig-RiR) model [12] introduces a nested RWKV structure with Outer and Inner blocks and a Zigzag-WKV attention mechanism to capture both global and local features while preserving spatial continuity. These innovations demonstrate RWKV’s ability to capture long-range dependencies, retain fine anatomical features, and scale to high-resolution images. However, systematic exploration of hybrid architectures combining RWKV with local feature extractors remains limited, leaving opportunities to balance local detail, global context, and computational efficiency in medical image segmentation.
III Methodology: RWKV-UNet
III-A Overall Architecture
@@
Figure 2: Overall architecture of the proposed RWKV-UNet. (a) the encoder with four stages constructed by stacked LP blocks and stacked GLSP blocks, (b) Cross-Channel Mix (CCM) Module for multi-scale fusion, (c) the decoder with four stages, (d) the Local Perception (LP) block, (e) the RWKV-based Global-Local Spatial Perception (GLSP) block, (f) the decoder block constructed by a point-convolution layer and a $9\times 9$ DW-Conv layer, with a convolution and an upsampling operation.
@@
The overall architecture of the proposed RWKV-UNet is presented in Fig. 2. RWKV-UNet consists of an encoder with stacked LP and GLSP blocks, a decoder, and skip connections with a Cross-Channel Mix (CCM) Module.
III-B Effective Encoder in RWKV-UNet
Global-Local Spatial Perception (GLSP) Block
@@ III-B Effective Encoder in RWKV-UNet
We employ the RWKV-Spatial-Mix module and Depth-Wise Convolution (DW-Conv) with a skip connection to combine global and local dependency. The process of expanding and reducing dimensions enhances feature representation by capturing richer details while preventing information bottlenecks. Given a feature map $X\in\mathbb{R}^{C_{\mathrm{in}}\times H\times W}$, the processing flow through the GLSP module is as follows: Normalize and project the input $X$ to an intermediate dimension $C_{mid}$ using a $1\times 1$ Convolution layer:
@@
Divide the feature map into patches of size $1\times 1$. Then the Spatial Mix in VRWKV is applied:
@@
Convert the feature sequence back into a 2D feature map $I_{4}=\text{Folding}(I_{3})\in\mathbb{R}^{C_{\mathrm{mid}}\times H\times W}$, and use a $5\times 5$ DW-Conv layer for local feature aggregation:
@@
Local Perception (LP) Block
An LP block contains a point convolution layer, a DW-Conv layer and a point convolution layer with local and global residual skips, removing the spatial mix as well as the unfolding and folding processes from the GLSP.
Effective Encoder with Stacked LPs and GLSPs
We construct the RWKV-UNet encoder by stacking the LP and GLSP blocks. The encoder comprises a stem stage followed by four stages. The first blocks in stages I, II, III and IV do not have residual connections and use a DW-Conv with a stride of 2 to achieve downsampling. After the first block, Stages I and II are composed of stacked LP blocks without Spatial Mix. In Stages III and IV, a series of GLSP blocks are stacked following the first block.
Scale Up
As shown in Table I, we implement three different sizes of encoders with different numbers of LP and GLSP blocks in each stage (depths), different embedding dimensions and expansion ratios for $C_{\mathrm{mid}}$ : Encoder-Tiny (Enc-T), Encoder-Small (Enc-S), and Encoder-Base (Enc-B), allowing flexibility to scale up the model according to different resource constraints and application needs.
Pre-training Manner
@@
TABLE I: RWKV-UNet Encoder Configurations Across Stages
@@
Parameters 2.94M 9.42M 16.69M
FLOPs 1.83G 4.73G 8.96G
TABLE II: Comparison results of models with different attention mechanisms on Synapse dataset.
@@
Multi-Head Self-Attention [2] 29.32 77.03 18.55G
Focused Linear Attention [44] 27.66 76.89 11.18G
Bi-Mamba [7] 29.10 78.13 14.53G
RWKV Spatial-Mix (ours) [10] 25.01 78.14 11.11G
@@
We conduct experiments by replacing the spatial mix in the encoder of RWKV-UNet with other attention mechanisms, such as self-attention [2], focused linear attention [44], and Bi-Mamba used in Vision Mamba [7], as well as by removing this module entirely. The total training epochs are 100. The results shown in Table II indicate that the model with Spatial-Mix achieves the best DSC with the lowest computational cost. Comparison visualizations of effective receptive fields of the last layer output using different attention mechanisms in the GLSP module are shown in Fig. 3.
(a) No Attention
(b) MHSA
(c) Bi-Mamba
(d) Spatial Mix
Figure 3: Comparison visualization of effective receptive fields of the last layer output using different attention mechanisms in the GLSP module.
III-C Large Kernel-based Decoder Design
We design our CNN-based decoder block with a point-convolution layer and a $9\times 9$ DW-Conv layer, following a point operation layer and an upsampling operation. Consider the input feature map $X\in\mathbb{R}^{C_{\text{in }}\times H\times W}$; in depthwise convolution, the number of input channels is $C_{\mathrm{in}}$, and the second point convolution maps the $C_{\text{in }}$ channels to $C_{\text{out }}$.
III-D Cross-Channel Mix for Multi-scale Fusion